OpenStudy (anonymous):

what is 7n-5=11n-17 ?

OpenStudy (mathhelp346):

what you want to do first is get n by it self

OpenStudy (mathhelp346):

so add 5 to -17 and you will get 7n=11n-12

OpenStudy (mathhelp346):

then you should subtract 11n from 7n and you will get -4n=-12

OpenStudy (mathhelp346):

do you understand so far?

OpenStudy (mathhelp346):

er..ok

OpenStudy (mathhelp346):

anyways, you multiple both sides by -1/4 in order to get n=-3

OpenStudy (anonymous):

The answer is 3. You have to move the 10 to right side of the equation which makes it positive so it is 11+10 on the right side. Then you divide both sides by 7 to isolate the n on the left side. 11+10 is 21 and 21 divided by 7 equals 3. :) 7n-10=11 7n=11+10 7n=21 7n/7=21/7 n=3

OpenStudy (mathhelp346):

*n=3

OpenStudy (anonymous):

yes